a 4 ohm resistance wire is doubled on it  calculate the new resistance of the wire.
for the above question, how to consider the question is for doubled the diameter or doubled the length.
Some says, the doubled on it means, increase the length.
Some says, the doubled on it means, diameter increased.
Now the yagya statement is
Now if you are put a resistance on the 4ohm ...then both will come in parallel and the new resistance will be 2 ohm 
explain in simple form
 
 

Grade:12

2 Answers

Saurabh Kumar
9 years ago
R = pL/A ---> Weather area of cross section or length chages, Volume remains same..
So define your formula in terms of Voulme...and proceed according to the requirement...
Yash Bhartiya
9 years ago
When the length of a conductor is varied , then its area also changes e.g.,when you stretch a wire ,it become thin . So, the length of the conductor changes from L1 to L2 and the area changes from A1 to A2
during this transformation , the volume of the conductor remain the same . So we have 
A1L1= A2L2
or A2= A1L1÷L2
Initially resistance of conductor 
R=pL1/A1
New resistance = R’=pL2/A2
now you can do by yourself